OpenStudy (anonymous):

Can someone please show me how you do this? Like the steps, because I'm kinda confused....

OpenStudy (anonymous):

Doing this using PEMDAS a = -3, b = 6, c = 3, and d = -5 |ac-d| |b-2|

OpenStudy (anonymous):

PEMDAS = Order of Operations (parenthesis, exponents, multiplication, division, addition, subtraction.)

OpenStudy (anonymous):

-3x3--5= -3x3+5

OpenStudy (anonymous):

For the first one, does that make sense?

OpenStudy (anonymous):

|-3+3+5| =5 |6-2| =4

OpenStudy (anonymous):

|ac-d| (over) |b-2| It's a fraction.

OpenStudy (anonymous):

All you pretty much do is plug in the number for the letter. ohhhhh... I see well then you will divide. 5\4 and get 1.25?

OpenStudy (anonymous):

im not convinced that its -3+3. Im pretty sure its -3 times by 3.

OpenStudy (anonymous):

well if it is -3(3)+5 then you answer will be -4/4 and you would get -1

OpenStudy (anonymous):

This is what I'm getting... what am I doing wrong?

OpenStudy (anonymous):

|-3*6-5|/|6-2| |-18-5|/|6-2| |-23|/|4| =-5.75

OpenStudy (anonymous):

you have 6 instead of 3. c=3

OpenStudy (anonymous):

OH, wow... thanks! Lol

OpenStudy (anonymous):

yep. :P

OpenStudy (anonymous):

|-3*3-5|/|6-2| |-9-5|/|6-2| |-14|/|4| =-3.5

OpenStudy (anonymous):

absolute value makes -14 positive :)

OpenStudy (anonymous):

a = -3, b = 6, c = 3, and d = -5 |ac-d| |b-2| |-3*3--5| = |-3*3+5| = |-9+5| = |-4| = 4 |6-2| = 4 4/4 = 1

jimthompson5910 (jim_thompson5910):

|ac - d|/|b-2| |(-3)(3) - (-5)|/|6-2| |-9 + 5|/|6-2| |-4|/|4| 4/4 1

jimthompson5910 (jim_thompson5910):

So this shows you that |ac - d|/|b-2| = 1 when a = -3, b = 6, c = 3 and d = -5